Use the absorption and fluorescence spectra of bacteriochlorophyll a dissolved in toluene to estimate:

(a) the Stokes shift of the fluorescence spectrum

(b) the frequency (in cm−1) of the vibrational mode that is displaced (stretched) by the optical transition from the ground state to the excited state. Compare the estimates you get from the vibrational structure of the absorption and fluorescence spectra. What part of the bacteriochlorophyll macrocycle is likely to be changing structure?Explain.

(c) the energy gaps in cm−1 between the lowest energy excited state and the next two higher excited states
Answer #1

Solution:

The literature source from royal society of chemistry is used as base reference for answering as below,

(a) The stokes shift of bacteriochlorophyll by fluorescence spectrum is,

B2 BUV737 Qdot 705 700 900 600 Wavelength (nm)

(b) The difference in spectrum falls under the range705-760 nm for the bacteriochlorophyll in the presence of solvent Toluene.

(c) \Delta E=h\Delta v=6.626*10^{-34}*\frac{1}{760nm-705nm}=1.205*10^{-26} J